June 15th

Added by Kyberd, Paul about 3 years ago

Report from analysis group

  • FC1 field mapping: Data taking is now completed in solenoid and flip mode, and the magnet remembers
    training in both operating modes.
  • Paper in progress: "Pion contamination in the Step I beamline," status will be updated at CM39
  • Paper in progress: "EMR commissioning," status will be updated at CM39

Focus Coils - good news.

After dismantling the field mapping equipment this morning, FC#1 immediately ramped up to 186.5A
in Flip mode before quenching, thus showing it remembers its training after mode changes.
We’ve subsequently made the snap decision to warm up over the weekend and get FC#2 cooling as soon as possible.

Also, FC#2 was leak checked today and showed not to have deteriorated significantly as a result of
the recent reparations at Tesla.

June 8th

Added by Kyberd, Paul about 3 years ago

Hall Summary, 4th June 2014

• Decay solenoid: new PSU in position, connections will start on Monday to allow start of commissioning – manufacturers arrive on 23rd June for final phase
• D2 PSU, which had to be moved, will also be reconnected Monday
• RR2 making good progress, concentrating on power distribution boards and cabling thereto, error with battery frame location will be corrected next Tuesday
• FC1 field map in flip mode complete, solenoid map in progress
• FC2 is being pumped down OK
• Vacuum hardware installation begins on Monday
• Downstream SS baseplate ready to be collected and modified by outside company (for compatibility with PRY)
• West mezzanine load test successful, deflection with design limits. Compressor frames now mostly built
• R78 target off for a couple of days, Paul Hodgson informed

June 1st

Added by Kyberd, Paul about 3 years ago

Hall Summary, 21st May 2014

- delayed because of problems with the news site

• Decay solenoid: successfully powered to 870 amps, QD tests complete, PSU will be swapped over next two weeks, then commissioned by manufacturers
• Automated proton absorber installation complete and tested and DSA closed
• D2 now back in position
• PRY bases delivered this week
• Distribution board and trunking installation making good progress in RR2
• Additional electrical technician employed w.e.f May 19th
• CERN field mapping gear is at RAL
• FC1 mapping begins 3rd June
• FC2 delivery (take two) on 23rd May
• Problem with target chiller now fixed
• Raised frame target run negated by above, will be repeated
• West mezzanine load test scheduled for May 30th, then wiring and compressor placement can go ahead
• European Cryogenics Group tour of R5.2 on 29th May
• Activation run on 27-29th June

Target news

R78 test S1.8 has exceeded 3.2M actuations and is performing as expected.

May 12th

Added by Kyberd, Paul about 3 years ago

The first Solenoid is in the Hall

The first solenoid was moved into the MICE hall last week.
images in the Photo Album

see also

Spokesman's update on Recent reviews of MICE

The technical successes across the project, the substantial improvements in organisation and the beginnings of effective preparation for operations at Step IV were all recognised and commended at last week’s reviews. As I write, the upstream solenoid, with tracker inside, is in the Hall on its stand (see the MICE photo repository), the downstream solenoid is in R9 awaiting the installation of the second tracker and the focus coil is being prepared for field mapping. We’re on our metal to deliver Step IV and, after all the years of preparation, the pressure to deliver is welcome.
The RLSR Panel and MPB discussed the programme needed for MICE to be scientifically and technically successful. In the MPB’s words (see the MPB site on MICEmine):
“MICE has 3 strategic technical and scientific goals. Demonstrate:
1) physical understanding of ionisation cooling with re-acceleration
2) technology of components integrated into an experimental arrangement
3) a working Neutrino Factory module
Goals 1 and 2 need Step V. Only goal 3 requires Step VI.”

We are all aware of the significant pressures on budgets across the collaboration and of the need to make best use of the tax-payers’ resources that we have been granted. In view of the budget pressures and the fact that Step V will allow the critical technology to be demonstrated and cooling with reacceleration to be measured, the MPB concluded that it did “not support proceeding to Step VI”.
The MPB also recommended that MICE “…explore all options for advancing the milestone for completing the installation of Step V in 2017.” This recognises the scientific and technical importance of Step V. Such an expedited Step V programme will reduce the total project cost (TPC). Clearly the continued, sustained and stable support of the funding agencies is vital for us to be able to deliver Step V on this expedited timescale. Initial steps towards the necessary negotiations with the funding agencies are in hand and Roy has initiated a review of the schedule.
Preparations for Step V are going well: the coupling coil remembers its training; the single-cavity test is being prepared; the RF-power amplifier refurbishment work is going well; the second LH2 absorber is on its way from Osaka; and the second focus coil will be returned to RAL from TESLA later this month.
Step IV is critical; only at Step IV shall we be able to study the cooling properties of LH2 and LiH, the materials that have been proposed for use in the cooling channel at the Neutrino Factory and Muon Collider. With continued hard work, attention to detail and focus we can bring the Step IV build to a successful conclusion, paving the way for success in Step IV operations and laying the foundations for the critical demonstration of ionization cooling with reacceleration at Step V.

Summary of the hall meeting

• Decay solenoid – compressor shaft seal changed again, cleaning system this week, hope to begin cooldown again on Friday 9th May
• Activation run likely in June – ISIS will give one week’s notice
• SS2 now in R5.2, aim to position in offline location on Friday 9th May
• SS1 is in R9
• Revised ODH plan for R5.2 to be complete for ISIS/MICE safety meeting on May 30th
• Dedicated services meeting on May 13th – priority will be layout of magnet cabling and compressor hoses
• Weekly phone meeting with DL established to expedite electrical work
• Survey work requested: PRY base fixings; tracker in SS1, FC1 field mapping gear
• Refit of MLCR underway
• Plan in hand to replace unreliable network switches
• Light curtain for PPS on order – one month delivery – ISIS are approving circuit mods
• Major parts of vacuum system now being installed in R5.2
• FC stable at 180amps for 48 hours, but quenched at end of run – not thought to be fault with magnet
• Mapping gear for FC1 about to be shipped from US
• FC2 due at RAL May 19th, FC1 still the baseline for Step IV

May 6th

Added by Kyberd, Paul about 3 years ago

Hall news

Today SS1 is arriving in R9 and SS2 is being moved to the hall

Target news

The development of the improved S series target is proceeding well with S1.8 currently
under test in R78. The goal is to have a high reliability target with a lifetime of
better than 5 million actuations (maybe 10!), the old T series has a lifetime of
1-2 million actuations. The current best is 10 million actuations for the S series
and the tests will continue to ensure the reproducibility of the target refits.

RF news from Fermilab

Progress toward operation of the first MICE production cavity at Fermilab

The first pair of RF couplers built at LBNL arrived at Fermilab and were assembled,
leak checked, test-fit on the cavity and adjusted Apr 16-18. Fine adjustment and more
detailed measurements with RF pickups were carried out the next week.

Over the past week, more instrumentation has been installed on the cavity and the vacuum
vessel was prepared for transport while in the clean room at Lab-6. Couplers were removed
and the cavity and vessel were sealed with port covers and transparent cover plates.
The move to the MTA took place on Friday, May 2nd.

The vessel is now in the MTA clean room on its final stand and couplers, cover plates and
miscellaneous other parts have also been moved to the MTA hall.
Final assembly will proceed over the next 2-3 weeks with more instrumentation, tuner actuators,
couplers and cover plates to be installed. A thorough inspection of the interior surface of the
cavity will also be performed before it is sealed for operation. The vessel will then be lifted
onto the platform, connected to services (air, water, vacuum) and baked in its final position.
Initial commissioning is expected to start in time for the MICE RF Workshop scheduled for June 2-3 at Fermilab.

Yagmur (who provided the update) proposes a caption contest for a photo (taken during the move by
Moses Chung) shown here

April 15th

Added by Kyberd, Paul about 3 years ago

Hall news

• Decay solenoid – seal repair completed 8th April, cooldown either underway or imminent
• R78 target re-fit completed on 7th April, new bearings and shaft, running resumed
• Proton absorber under final test, install this shutdown
• SS2 move to R5.2 confirmed for 29th April
• Beamstopper should be fixed April 11th
• Request to survey online target during shutdown
• Bolt will be fitted to top of Potterax door (to solve PPS problems)
• PPD safety tour on 16th April, frantic tidy up in progress
• New MOM phone being purchased
• Rack room two will be ready for flooring 14th April
• TIARA/UMISS RF gear will be packed for DL April 16th
• Easter shutdown dates: 7th April – 6th May

MOM update

  • MICE ran on Sunday 6th April to test the new target double dip rate.
    The target performed as expected and generated no problems in ISIS. Everyone is happy.
  • The new DAQ was also tested with data and appeared to be working well too.
  • A few niggling issues (glitch with the PPS, repair needed to the beam stop, a walk
    in during the hall search) but all in all a very successful shift

Current Mom

From Monday 14th: Ray Gamet

April 7th

Added by Kyberd, Paul about 3 years ago

Hall Report

  • Decay solenoid – seal repair confirmed for Tuesday, 8th April
  • Vacuum system mock-up complete
  • Safety tour of R9 & R5.2 on April 16th
  • RR2 flooring and cabling in progress, on schedule for rack delivery in May
  • Tracker installation complete, site visit for solenoid movement 8th April

Current MOM

Adam Dobbs

March 31st

Added by Kyberd, Paul about 3 years ago

Hall Summary, 27-3-14

• Decay solenoid – fridge compressor replacement seal expected within two weeks
• A new problem with Linde chiller – controls and gas level
• Air conditioning in Hall fixed and available
• Ex. TIARA RF gear stripped and ready to ship to DL (for commissioning second TH116)
• RR2 cable tray work making good progress, then raised floor can be finalised
• FC#1 reached 188.5 amps in flip mode; inactive this week due to tracker installation, but will resume soon with extended stability testing
• Tracker installation in SS#2 should be complete this week; transport to R5.2 tentatively arranged for last week April
• PSU in DAQ rack now fixed
• R78 target running again
• West end crane back in use – minor defect with aux. hoist on East crane
• Conventional magnets exercised OK in readiness for April 6th run

Tracker Installation

The tracker installation is almost complete. Having ordered a few missing screws that we needed to complete things, we will be attaching the lightguides to the patch panel on Monday. We have been taking pictures throughout which will be posted very soon.


Also available in: Atom